OpenStudy (anonymous):

Rewrite each expression without absolute value bars. |sqrt2 - 5| I have the answer, but I have no idea how we get there or why. I would appreciate an explanation :)

jimthompson5910 (jim_thompson5910):

The problem is \[\Large \left|\sqrt{2}-5\right|\] correct?

OpenStudy (anonymous):

Yes.

jimthompson5910 (jim_thompson5910):

Since 2^2 = 4, this means that the square root of 2 is less than 5 So \[\Large \sqrt{2} < 5\] ---------------- This means that \[\Large \sqrt{2} < 5\] \[\Large \sqrt{2} - 5 < 0\]

jimthompson5910 (jim_thompson5910):

Now use the idea that |x| = x ... if x >= 0 or |x| = -x .... if x < 0